# Entanglement in quenched extended Su-Schrieffer-Heeger model with anomalous dynamical quantum phase transitions

### v1

@@ V Conclusion
We attempted to integrate several types of DQPT into the same framework, namely, the type-B anomalous DQPT, by comparing it with regular DQPTs in the extended SSH model with two chiral-symmetry-preserving next-nearest-neighbor hopping terms added to the original SSH model. The introduction of these hopping terms induces two additional phases on top of $W_{0}$ and $W_{1}$, namely $W_{2}$ and $W_{-1}$, in equilibrium. We investigated the dynamical phase diagram by counting the number of critical momenta $n_{k^{*}}$ of the corresponding quench, and defined the type-B anomalous DQPTs as $n_{k^{*}}$ bigger than the absolute winding number difference between the pre- and postquenched phases. Under this formulation, type-B anomalous DQPT incorporates the possibility of DQPTs occur for quenches within the same phase (both with and without crossing equilibrium phase boundary), and quenches that involve multiple $n_{k^{*}}$ and thus critical time scales. On top of that, we elucidated the extra critical momenta always come in a pair due to the fact that they are complex conjugate pair of each other. The trajectories of the dynamical vector $\vec{r}_{k}$ helps to distinguish the nature of the regular critical momentum and the extra critical momenta, where the latter pair stays in the same loop of trajectory instead of two separate loops.
@@
Furthermore, we classified the quench scenarios by the equilibrium entanglement entropy phase diagram, where $W_{2}$ and $W_{-1}$ are the more non-local (overall higher EE) phases forming class 2, while $W_{0}$ and $W_{1}$ are the less non-local (lower EE) phases forming class 1. As such, there are 5 scenarios of entropy dynamics quench, namely, from class 1 to class 2, and vice versa, from class 2 to itself with/without changing winding number, and from class 1 to itself. We discussed the time evolution of the EE near the first DQPT regarding these 5 different situations. On one hand, if we quench deep inside the phase, DQPT occurs near the local maximum of EE for class 1 to class 2 and from class 2 to itself with $|\Delta W|=n_{k^{*}}$, regardless of the size of the subsystem. In these quench cases, the deeper the quench to the phase, the more alignment between the time EE is maximized and the first critical time. The cause of this maximum comes from the level crossings occurred among the correlation matrix spectrum, giving the highest momentum uncertainty during the reorganization of correlation during DQPTs.
@@
On the other hand, with an appropriate subsystem choice, the DQPT occurs near the local minimum of the entanglement entropy for quenches from class 2 to class 1 and class 2 to itself with $|\Delta W|<n_{k^{*}}$. The variation in time of the correlation matrix spectrum in this case is the least and the effective eigen-levels are farthest away from the Fermi level near the critical time. In addition, one might be able to realize a quench that attains lower entropy with a small subsystem near DQPT compared to the pre-quenched phase. A possible application is that one can utilize such property to perform measurement on just a small subsystem, yet still capable of capturing significant amount of information of the whole system comparing to its pre-quenched phase. For the quenches within class-1 phases, it is consistent with the original SSH model, where $W_{1}\rightarrow W_{0}$ and $W_{0}\rightarrow W_{1}$ attain minimum and maximum for subsystem $L_{A}=1$ in the vicinity of DQPT respectively. Our results show $1k^{*}$ and $3k^{*}$ regions have second and first maximum/minimum in EE near the first critical time respectively. We interpret the first minimum/maximum in EE near DQPT in $1k^{*}$ quench as a preparation before entering the $3k^{*}$ region.
@@
Our work may serve as an attempt to extend the study of DQPTs in topological models with more than one topological phases with distinct topological characters. The question of any physical quantities like order parameter, whether equilibrium or dynamical, can detect the anomalous regions studied in this work remains to be explored. Therefore, topological models that happen to display type-B anomalous dynamical regions in the dynamical phase diagram are of particular interest in the future research. A natural extension would be to generalize the definition of type-B anomalous DQPTs to other models of different topological invariants such as Chern number [55, 97] and models in higher dimensions. For example, for models without chiral symmetry where the winding number is not a good topological invariant [54]. It is also intriguing to explore the applicability of our results to generalization of unit cells instead of hopping range, which also exhibits type-B anomalous DQPTs [55]. On the other hand, the Loschmidt amplitude spectrum [33, 41] may also serve as a probe to study the physical effect of the extra critical momenta in topological models in the future work.
@@ Acknowledgements.
We acknowledge financial support from Research Grants Council of Hong Kong (Grant No. CityU 21304020), FCT through Grant No. UID/CTM/04540/2019, and City University of Hong Kong (Grant No. 9610438, 7005610, 9680320).
@@ Appendix A Type-B anomalous DQPTs in the SSH model with the next-next-nearest-neighbor hopping
The dynamical phase diagram is shown in Fig. 14(a) and Fig. 14(b) with prequenched phase as $W_{2}$ and the lower $W_{3}$ phase, respectively. Similar to the dynamical phase diagram in Fig. 2(b) and 2(c) for $t_{e}=t_{f}=0$, anomalous regions with extra pair(s) of $k^{*}$’s are found for quench $W_{2}\rightarrow W_{-1}$ and $W_{3}\rightarrow W_{3}$. We notice a couple of differences regarding the initial phase as the lower $W_{3}$ phase under the presence of the longer-range hoppings (compare Fig. 14(b) and Fig. 2(c)): There is a small anomalous region near the phase boundaries $t_{f}=-t_{e}-1$ and $t_{f}=t_{e}$ of $W_{2}$, and also $t_{f}=t_{e}$ of the upper $W_{3}$; a tiny anomalous region of $4k^{*}$’s besides the $2k^{*}$ region appears inside the lower $W_{3}$ phase near the multicritical point $(t_{e}^{c},t_{f}^{c})=(-0.5,-0.5)$. Nevertheless, the critical momenta pairs in $\vec{r}_{k}$ of the anomalous DQPTs behave the same as described in the main text, i.e. they stay in one loop as shown in Fig. 14(d) and 14(e). Additionally, the dynamical phase diagram for $t_{a}=t_{b}=-1,t_{c}=t_{d}=0$ from the $W_{1}$ phase is also shown in Fig. 14(c). For the quench to the $W_{0}$ phase, there are two anomalous regions, one with $3k^{*}$ and one with $5k^{*}$, observed. The $\vec{r}_{k}$ trajectory for a quench to the $5k^{*}$ region is depicted in Fig. 14(f). It clearly shows two pairs of extra $k^{*}$’s, with each pair staying in one loop, again validating the complex-conjugate-pair statement presented in the main text.
@@ A.1 Entanglement entropies
From Table 2, note that for a phase in Class 2 quenching to that in Class 1, a local minima of the EE can be achieved in the vicinity of DQPT with a suitable choice of the subsystem. The finding may also be generalized to other cases with quench from a higher to a lower classes, and this can be confirmed by Fig. 15(a), where the EE for quenches from $W_{3}$ to phases in all the classes, including $W_{1}$ (Class 1), $W_{-1}$ (Class 2), and $W_{3}$ (Class 3), are presented. It clearly shows the critical time occurs near the EE’s local minima for quench from $W_{3}$ to $W_{1}$ and $W_{-1}$. In the main text, we also find that the anomalous DQPT resulted from the quench from $W_{2}$ or $W_{-1}$ to itself exhibits a local minima in the EE with a suitable subsystem choice (see Table 2, Fig. 10 (b) and (c)). The quench for $W_{3}$ and $W_{-2}$ should also follow similar pattern and this is validated by the pink and purple curves in Fig. 15(a). The EE for the $W_{3}\rightarrow W_{3}$ quench exhibits a local minima near DQPT, regardless of $n_{k^{*}}=2$ or $n_{k^{*}}=4$. We have verified the above also holds for quench initially from the $W_{-2}$ phase with a subsystem of $L_{B}=3$ (not shown here).
@@
In the main text, we also find that local maxima in the EE can be attained around DQPT for deep quenches from Class 1 to Class 2 (see Fig. 9(b)), or from Class 2 to Class 2 with a regular DQPT (see Fig. 10(a)). One infers that deep quenches from a lower to a higher class or between phases in the same class with regular DQPTs will also achieve a local maxima in EE around the critical time. In order words, we expect deep quenches from $W_{-1/0/1/2}$ to $W_{-2/3}$ or $W_{-2}/W_{3}$ to $W_{3}/W_{-2}$ shows local maxima in EE around DQPTs. Figure 15(b) shows a deep phase quench of $W_{-1}\rightarrow W_{3}$, where the darker the color of the curve is, the deeper the final parameter is inside the phase. We indeed observe maxima of EE around DQPTs. In addition, the deeper into the $W_{3}$ phase, the more the EE maxima and the critical time align. This is consistent with what we have obtained from the inset of Fig. 10(a).
@@
Figure 16: Time evolution of the correlation matrix spectra for an $L=80$ system with (a-d) $t_{a}=t_{b}=t_{c}=t_{d}=-1$ and (e-g) $t_{a}=t_{b}=-1,t_{c}=t_{d}=0$ (see Fig. 14(a), 14(b) and 14(c) for the dynamical phase diagrams respectively). The quench cases are (a) $(t_{e}^{i}=-1,t_{f}^{i}=2.5)\rightarrow(t_{e}^{f}=1.9,t_{f}^{f}=0.9)$, $W_{2}\rightarrow W_{-1}$ inside $3k^{*}$ region; (b) $(t_{e}^{i}=-1,t_{f}^{i}=2.5)\rightarrow(t_{e}^{f}=3.9,t_{f}^{f}=0.8)$, $W_{2}\rightarrow W_{-1}$ inside $5k^{*}$ region; (c) $(t_{e}^{i}=-1.8,t_{f}^{i}=-3.6)\rightarrow(t_{e}^{f}=-0.5,t_{f}^{f}=-2.1)$, $W_{3}^{-}\rightarrow W_{3}^{-}$ without DQPT, (d) $(t_{e}^{i}=-1.8,t_{f}^{i}=-3.6)\rightarrow(t_{e}^{f}=-0.5,t_{f}^{f}=-1.4)$, $W_{3}^{-}\rightarrow W_{3}^{-}$ with DQPT; (e) $(t_{e}^{i}=0.5,t_{f}^{i}=-1)\rightarrow(t_{e}^{f}=-0.5,t_{f}^{f}=-0.1)$, $W_{-1}\rightarrow W_{0}$ inside $1k^{*}$ region; (f) $(t_{e}^{i}=0.5,t_{f}^{i}=-1)\rightarrow(t_{e}^{f}=-1.05,t_{f}^{f}=0.5)$, $W_{-1}\rightarrow W_{0}$ inside $3k^{*}$ region; (g) $(t_{e}^{i}=0.5,t_{f}^{i}=-1)\rightarrow(t_{e}^{f}=-1.6,t_{f}^{f}=1.1)$, $W_{-1}\rightarrow W_{0}$ inside $5k^{*}$ region. Top, middle and bottom rows represent subsystems $L_{A}=2,4,40$ respectively.
@@ A.2 Correlation matrix spectra
We adopt the notation for the upper/lower $W_{3}$ phases as $W_{3}^{+/-}$ in this section. Figure 16 compares the time evolution of the correlation matrix spectrum between quenches to type-B anomalous and regular regions. The eigen-levels for $W_{2}\rightarrow W_{-1}$ with anomalous DQPT (Fig. 16(b)) is more entangled and concentrated around the Fermi level, and at the same time the quench triggers slightly more level crossings as compared to the regular case (Fig. 16(a)). This is similar to the observations for the $W_{1}\rightarrow W_{0}$ case (see Fig. 12(a) and 12(b)) in the main text. Note however that for the half-block’s correlation matrix spectrum, level crossing at the Fermi level are found in both the regular and the type-B anomalous cases (the bottom panel of Fig. 16(a) and 16(b)), unlike the case for $W_{1}\rightarrow W_{0}$ where quenches to the regular region (Fig. 12(a)) do not trigger crossings for a reasonably long time. The difference may arise from the fact that quench between two topological phases is involved for $W_{2}\rightarrow W_{-1}$, as compared to the quench between a topological phase and a trivial phase for $W_{1}\rightarrow W_{0}.$ On the other hand, if one considers the quench from a topological $W_{-1}$ phase to a trivial $W_{0}$ phase in the further extended SSH model, as shown in Fig. 16(e-g), the distinctive features between the anomalous and the regular DQPTs are found to be consistent with those presented in Sec. IV. Namely, persistent double degeneracy at the Fermi level with the degenerate levels gapped from the rest in the regular case ($1k^{*}$), and the levels tend to mix and cross for both the $3k^{*}$ and $5k^{*}$ anomalous cases.
@@
The quenches within the $W_{3}^{-}$ phases shown in Fig. 16(c-d) also resemble the dynamics of the correlation matrix spectrum for quenches within the $W_{2}^{-}$ phase shown in Fig. 12(c-d). For the subsystem of $L_{A}=2,4$, the spectrum has more crossings in the regular case as compared to the anomalous case. Furthermore, the eigen-levels around the Fermi level for the half-block subsystem are gapped from the rest in the regular case, whereas the eigen-levels away from the Fermi level are more involved in the anomalous case. We also note that there is a persistent double degeneracy at the Fermi level in the half-block correlation matrix spectrum for the quench from $W_{3}$, which is absent in the case of quench from $W_{2}$ in the main text. The subsystem recovers the six-fold degeneracy temporarily at the time where level crossing at the Fermi level occurs. The persistent double degeneracy in the half-block subsystems is in fact guaranteed for quenches from a phase with odd winding number. Note that the eigen-levels whose eigenvalues are neither 0 nor 1 are doubly degenerate. Since initially the phase has $2W$ edge states, the phases with odd winding number must have a pair of eigen-levels left which does not belong to either upper or lower branch of the spectrum, given the spectrum are symmetric along the Fermi level due to the presence of particle-hole symmetry. That one pair has to stay at the Fermi level, leading to a persistent double degeneracy happened in quenches from odd-winding-number phases.

### v2

@@ V Conclusion
We attempted to integrate several types of DQPTs into the same framework, namely, the type-B anomalous DQPTs, by comparing it with regular DQPTs in the extended SSH model with two chiral-symmetry-preserving next-nearest-neighbor hopping terms added to the original SSH model. The introduction of these hopping terms induces two additional phases on top of $W_{0}$ and $W_{1}$, namely $W_{2}$ and $W_{-1}$, in equilibrium. We investigated the dynamical phase diagram by counting the number of critical momenta $n_{k^{*}}$ of the corresponding quench, and defined the type-B anomalous DQPTs as $n_{k^{*}}$ bigger than the absolute winding number difference between the pre- and postquenched phases. Under this formulation, type-B anomalous DQPTs incorporate the possibility of DQPTs occuring for quenches within the same phase (both with and without crossing equilibrium phase boundary), and quenches that involve multiple $n_{k^{*}}$ and thus critical time scales. On top of that, we elucidated the extra critical momenta always come in a pair due to the fact that they are complex conjugate pair of each other. The trajectories of the dynamical vector $\vec{r}_{k}$ helps to distinguish the nature of the regular critical momenta and the extra critical momenta, where the latter pair stays in the same loop of trajectory instead of two separate loops.
@@
Furthermore, we classified the quench scenarios by the equilibrium entanglement entropy phase diagram, where $W_{2}$ and $W_{-1}$ are the more non-local (overall higher EE) phases forming class 2, while $W_{0}$ and $W_{1}$ are the less non-local (lower EE) phases forming class 1. As such, there are six scenarios of entropy dynamics quench, namely, from class 1 to class 2, and vice versa, from class 2 to itself with/without changing winding number, and from class 1 to itself with $1k^{*}$/ $3k^{*}$. We discussed the time evolution of the EE near the first DQPT regarding these six different situations. On one hand, if we quench deep inside the phase, DQPT occurs near the local maximum of EE for class 1 to class 2 and from class 2 to itself with $n_{k^{*}}=|\Delta W|$, regardless of the size of the subsystem. In these quench cases, the deeper the quench to the phase, the more alignment between the time EE is maximized and the first critical time. The cause of this maximum comes from the level crossings occurred among the correlation matrix spectrum, giving the highest momentum uncertainty during the reorganization of correlation during DQPTs.
@@
On the other hand, with an appropriate subsystem choice, the DQPT occurs near the local minimum of the entanglement entropy for quenches from class 2 to class 1 and class 2 to itself with $n_{k^{*}}>|\Delta W|$. The variation in time of the correlation matrix spectrum in this case is the least and the effective eigen-levels are farthest away from the Fermi level near the critical time. In addition, one might be able to realize a quench that attains lower entropy with a small subsystem near DQPT compared to the pre-quenched phase. A possible application is that one can utilize such property to perform measurement on just a small subsystem, yet still capable of capturing significant amount of information of the whole system comparing to its pre-quenched phase. For the quenches within class-1 phases with $3k^{*}$, it is consistent with the original SSH model, where $W_{1}\rightarrow W_{0}$ and $W_{0}\rightarrow W_{1}$ attain minimum and maximum for subsystem $L_{A}=1$ in the vicinity of DQPT respectively. Our results show DQPT occurs around the second and the first minimum of the EE time evolution in the $1k^{*}$ and $3k^{*}$ regions, respectively. We interpret the first minimum in EE in the $1k^{*}$ region as a preparation for the DQPT that emerges when entering the $3k^{*}$ region.
@@
Our work may serve as an attempt to extend the study of DQPTs in topological models with more than one topological phases with distinct topological characters. The question of any physical quantities like the order parameter, whether equilibrium or dynamical, can detect the anomalous regions studied in this work remains to be explored. Topological models that happen to display type-B anomalous dynamical regions in the dynamical phase diagram are of particular interest in the future research. A natural extension would be to generalize the definition of type-B anomalous DQPTs to other models of different topological invariants such as Chern number [55, 97] and models in higher dimensions. For example, for models without chiral symmetry where the winding number is not a good topological invariant [54]. It is also intriguing to explore the applicability of our results to generalization of unit cells instead of hopping range, which also exhibits type-B anomalous DQPTs [55]. Besides, investigating the type-B anomalous DQPTs under the influence of non-Hermitian couplings in the SSH model [102] also presents an interesting research direction. On the other hand, the Loschmidt amplitude spectrum [33, 41] may also serve as a probe to study the physical effect of the extra critical momenta in topological models in the future work.
@@ Acknowledgements.
We acknowledge financial support from Research Grants Council of Hong Kong (Grant No. CityU 21304020), FCT through Grant No. UID/CTM/04540/2019, and City University of Hong Kong (Grant No. 9610438, No. 7006018, and No. 9680320).
@@ Appendix A Type-B anomalous DQPTs in the SSH model with the next-next-nearest-neighbor hopping
The dynamical phase diagram for $t_{a}=t_{b}=t_{c}=t_{d}=-1$ is shown in Figs. 14(a) and 14(b) with prequenched phase as $W_{2}$ and the lower $W_{3}$ phase, respectively. Similar to the dynamical phase diagram in Figs. 2(b) and 2(c) for $t_{e}=t_{f}=0$, anomalous regions with extra pair(s) of $k^{*}$’s are found for quench $W_{2}\rightarrow W_{-1}$ and $W_{3}\rightarrow W_{3}$. We notice a couple of differences regarding the initial phase as the lower $W_{3}$ phase under the presence of the longer-range hoppings (compare Fig. 14(b) and Fig. 2(c)): There is a small anomalous region near the phase boundaries $t_{f}=-t_{e}-1$ and $t_{f}=t_{e}$ of $W_{2}$, and also $t_{f}=t_{e}$ of the upper $W_{3}$; a tiny anomalous region of $4k^{*}$’s besides the $2k^{*}$ region appears inside the lower $W_{3}$ phase near the multicritical point $(t_{e}^{c},t_{f}^{c})=(-0.5,-0.5)$. Nevertheless, the critical momenta pairs in $\vec{r}_{k}$ of the anomalous DQPTs behave the same as described in the main text, i.e. they stay in one loop as shown in Figs. 14(d) and 14(e). Additionally, the dynamical phase diagram for $t_{a}=t_{b}=-1,t_{c}=t_{d}=0$ from the $W_{1}$ phase is also shown in Fig. 14(c). For the quench to the $W_{0}$ phase, there are two anomalous regions, one with $3k^{*}$ and one with $5k^{*}$, observed. The $\vec{r}_{k}$ trajectory for a quench to the $5k^{*}$ region is depicted in Fig. 14(f). It clearly shows two pairs of extra $k^{*}$’s, with each pair staying in one loop, again validating the complex-conjugate-pair statement presented in the main text.
@@ A.1 Entanglement entropies
From Table 2, note that for a phase in Class 2 quenching to that in Class 1, a local minimum of the EE can be achieved in the vicinity of DQPT with a suitable choice of the subsystem. The finding may also be generalized to other cases with quench from a higher to a lower classes, and this can be confirmed by Fig. 15(a), where the EE for quenches from $W_{3}$ to phases in all the classes, including $W_{1}$ (Class 1), $W_{-1}$ (Class 2), and $W_{3}$ (Class 3), are presented. It clearly shows the critical time occurs near the EE’s local minima for quench from $W_{3}$ to $W_{1}$ and $W_{-1}$. In the main text, we also find that the anomalous DQPT resulted from the quench from $W_{2}$ or $W_{-1}$ to itself exhibits a local minimum in the EE with a suitable subsystem choice (see Table 2, Figs. 10(b) and 10(c)). The quench for $W_{3}$ and $W_{-2}$ should also follow similar pattern and this is validated by the pink and purple curves in Fig. 15(a). The EE for the $W_{3}\rightarrow W_{3}$ quench exhibits a local minimum near DQPT, regardless of $n_{k^{*}}=2$ or $n_{k^{*}}=4$. We have verified the above also holds for quench initially from the $W_{-2}$ phase with a subsystem of $L_{B}=3$ (not shown here).
@@
In the main text, we also find that local maxima in the EE can be attained around DQPT for deep quenches from Class 1 to Class 2 (see Fig. 9(b)), or from Class 2 to Class 2 with a regular DQPT (see Fig. 10(a)). One infers that deep quenches from a lower to a higher class or between phases in the same class with regular DQPTs will also achieve a local maximum in EE around the critical time. In order words, we expect deep quenches from $W_{-1/0/1/2}$ to $W_{-2/3}$ or $W_{-2}/W_{3}$ to $W_{3}/W_{-2}$ to show local maxima in EE around DQPTs. Figure 15(b) shows a deep phase quench of $W_{-1}\rightarrow W_{3}$, where the darker the color of the curve is, the deeper the final parameter is inside the phase. We indeed observe maxima of EE around DQPTs. In addition, the deeper into the $W_{3}$ phase, the more the EE maxima and the critical time align. This is consistent with what we have obtained from the inset of Fig. 10(a).
@@
Figure 16: Time evolution of the correlation matrix spectra for an $L=80$ system with (a-d) $t_{a}=t_{b}=t_{c}=t_{d}=-1$ and (e-g) $t_{a}=t_{b}=-1,t_{c}=t_{d}=0$ (see Fig. 14(a), 14(b) and 14(c) for the dynamical phase diagrams respectively). The quench cases are (a) $(t_{e}^{i}=-1,t_{f}^{i}=2.5)\rightarrow(t_{e}^{f}=1.9,t_{f}^{f}=0.9)$, $W_{2}\rightarrow W_{-1}$ inside $3k^{*}$ region; (b) $(t_{e}^{i}=-1,t_{f}^{i}=2.5)\rightarrow(t_{e}^{f}=3.9,t_{f}^{f}=-0.8)$, $W_{2}\rightarrow W_{-1}$ inside $5k^{*}$ region; (c) $(t_{e}^{i}=-1.8,t_{f}^{i}=-3.6)\rightarrow(t_{e}^{f}=-0.5,t_{f}^{f}=-2.1)$, $W_{3}^{-}\rightarrow W_{3}^{-}$ without DQPT, (d) $(t_{e}^{i}=-1.8,t_{f}^{i}=-3.6)\rightarrow(t_{e}^{f}=-0.5,t_{f}^{f}=-1.4)$, $W_{3}^{-}\rightarrow W_{3}^{-}$ with DQPT; (e) $(t_{e}^{i}=0.5,t_{f}^{i}=-1)\rightarrow(t_{e}^{f}=-0.5,t_{f}^{f}=-0.1)$, $W_{1}\rightarrow W_{0}$ inside $1k^{*}$ region; (f) $(t_{e}^{i}=0.5,t_{f}^{i}=-1)\rightarrow(t_{e}^{f}=-1.05,t_{f}^{f}=0.5)$, $W_{1}\rightarrow W_{0}$ inside $3k^{*}$ region; (g) $(t_{e}^{i}=0.5,t_{f}^{i}=-1)\rightarrow(t_{e}^{f}=-1.6,t_{f}^{f}=1.1)$, $W_{1}\rightarrow W_{0}$ inside $5k^{*}$ region. Top, middle and bottom rows represent subsystems $L_{A}=2,4,40$ respectively.
@@ A.2 Correlation matrix spectra
We adopt the notation for the upper/lower $W_{3}$ phases as $W_{3}^{+/-}$ in this section. Figure 16 compares the time evolution of the correlation matrix spectrum between quenches to type-B anomalous and regular regions. The initial quench parameters in Figs. 16(a)-16(b), 16(c)-16(d), and 16(e)-16(g) are the same as that in the dynamical phase diagram shown in Figs. 14(a), 14(b), and 14(c), respectively. The eigen-levels for $W_{2}\rightarrow W_{-1}$ with anomalous DQPT (Fig. 16(b)) is more entangled and concentrated around the Fermi level, and at the same time the quench triggers slightly more level crossings as compared to the regular case (Fig. 16(a)). This is similar to the observations for the $W_{1}\rightarrow W_{0}$ case (see Figs. 12(a) and 12(b)) in the main text. Note however that for the half-block’s correlation matrix spectrum, level crossing at the Fermi level are found in both the regular and the type-B anomalous cases (the bottom panel of Figs. 16(a) and 16(b)), unlike the case for $W_{1}\rightarrow W_{0}$ where quenches to the regular region (Fig. 12(a)) do not trigger crossings for a reasonably long time. The difference may arise from the fact that quench between two topological phases is involved for $W_{2}\rightarrow W_{-1}$, as compared to the quench between a topological phase and a trivial phase for $W_{1}\rightarrow W_{0}.$ On the other hand, if one considers the quench from a topological $W_{1}$ phase to a trivial $W_{0}$ phase in the further extended SSH model, as shown in Figs. 16(e)-16(g), the distinctive features between the anomalous and the regular DQPTs are found to be consistent with those presented in Sec. IV. Namely, persistent double degeneracy at the Fermi level with the degenerate levels gapped from the rest in the regular case ($1k^{*}$), and the levels tend to mix and cross for both the $3k^{*}$ and $5k^{*}$ anomalous cases.
@@
The quenches within the $W_{3}^{-}$ phases shown in Figs. 16(c)-16(d) also resemble the dynamics of the correlation matrix spectrum for quenches within the $W_{2}^{-}$ phase shown in Figs. 12(c)-12(d). For the subsystem of $L_{A}=2,4$, the spectrum has more crossings in the regular case as compared to the anomalous case. Furthermore, the eigen-levels around the Fermi level for the half-block subsystem are gapped from the rest in the regular case, whereas the eigen-levels away from the Fermi level are more involved in the anomalous case. We also note that there is a persistent double degeneracy at the Fermi level in the half-block correlation matrix spectrum for the quench from $W_{3}$, which is absent in the case of quench from $W_{2}$ in the main text. The subsystem recovers the six-fold degeneracy temporarily at the time where level crossing at the Fermi level occurs. The persistent double degeneracy in the half-block subsystems is in fact guaranteed for quenches from a phase with odd winding number. Note that the eigen-levels whose eigenvalues are neither 0 nor 1 are doubly degenerate. Since initially the phase has $2W$ edge states, the phases with odd winding number must have a pair of eigen-levels left which does not belong to either upper or lower branch of the spectrum, given the spectrum are symmetric along the Fermi level due to the presence of particle-hole symmetry. That one pair has to stay at the Fermi level, leading to a persistent double degeneracy happened in quenches from odd-winding-number phases.